OpenStudy (anonymous):

the median of a set of data is ___ equal to the third quartile. a. sometimes b. always c. never

OpenStudy (cwrw238):

i'm a bit puzzled by this one, then statistics is not my strong point. Its certainly not b. always

OpenStudy (cwrw238):

the third quartile is i believe the median of the data above the median. I suppose if the data consists of say 10 items value 8 then its true lol! so I go for a sometimes

OpenStudy (amistre64):

median = Q2

OpenStudy (amistre64):

sometimes seems right to me as well 1 2 2 2 4 Q2 = 2, Q3 = 5*3/4 = 15/4 = 3... 4th postition = 2

OpenStudy (amistre64):

are there more precise ways to calculate the Qs? i only learneded this method

OpenStudy (cwrw238):

yea me too

OpenStudy (amistre64):

one oddity is that "quartile" has different meanings it can refer to the range of values between Q2 and Q3 .... or it can refer to the value of Q3

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

http://mathforum.org/library/drmath/view/60969.html im sure the solution is "sometimes" tho .... otherwise there would be no definitive answer :)